### v3

# Leibniz rule for wedge product in discrete exterior calculus on general polygonal meshes

Abstract
Discrete exterior calculus offers a coordinate–free discretization of exterior calculus especially suited for computations on curved spaces. In this work, we present a wedge product on 2–dimensional pseudomanifolds, whose faces are any polygons. We prove that this polygonal wedge product is compatible with the discrete exterior derivative in the sense that it satisfies the Leibniz product rule. We thus extend previously studied discretizations of wedge products from simplicial or quadrilateral meshes to surface meshes whose faces are arbitrary simple polygons. We also prove that our discrete wedge product corresponds to a cup product of cochains on 2–pseudomanifolds. By rigorously justifying our construction we add another piece to ever evolving discrete versions of exterior calculus.
keywords:
Discrete wedge product , Discrete exterior calculus , Polygonal wedge product , Polygonal cup product , Discrete differential forms , Polygonal exterior calculus.
MSC:
65N22 , 57Z20 , 57-08 , 53Z30
[…]
organization=Dept. of Applied Mathematics, Charles University, addressline=Ke Karlovu 3, city=Praha 2, postcode=121 12, country=Czech Republic
1 Introduction
The exterior calculus of differential forms, introduced by Élie Joseph Cartan (1869 – 1951), has become the foundation of modern differential geometry. It is a coordinate–free calculus and therefore greatly simplifies analysis and calculations on curved spaces of differential manifolds.
Discrete versions of exterior calculus on surface meshes have gained increased attention in the last few decades, see Hirani (2003); Arnold (2018); Desbrun et al. (2008); Ptackova and Velho (2021) and references therein. Discete exterior calculus (DEC) and finite element exterior calculus (FEEC) have found applications in computational fluid dynamics (Mohamed et al., 2016; Nitschke et al., 2017), computational electromagnetism (Zhang et al., 2021), and computer graphics (Desbrun et al., 2003; Ptackova and Velho, 2021). See also references in Schubel et al. (2024) for more applications.
In this paper, we prove that the wedge product on general polygonal meshes, first introduced in Ptackova and Velho (2017) and later used in applications in Ptackova and Velho (2021), satisfies the Leibniz product rule with discrete exterior derivative (coboundary operator). Furthermore, we show that it corresponds to a cup product of cochains on orientable 2–dimensional pseudomanifolds. We thus partially extend the findings of Schubel et al. (2024) from simplicial to general polygonal setting.
On smooth manifolds, the wedge product allows for building higher degree forms from lower degree ones. Similarly a cup product is a product of two cochains of arbitrary degree $p$ and $q$ that returns a cochain of degree $p+q$ located on $(p+q)$–dimensional cells. The cup product was introduced by J. W. Alexander, E. Čech, and H. Whitney (Whitney, 1957) in 1930’s and it became a well–studied notion in algebraic topology, mainly in the simplicial setting (Munkres, 1984; Fenn, 1983). Later, the cup product was extended also to $n$–cubes (Massey, 1991). In (Arnold, 2012) various cup products on simplicial and cubical $n$–pseudomanifolds are presented. Arnold (2012) also defines L2 cubical Whitney forms and a cubical cup product that fit together with the wedge product of cubical Whitney forms.
In (Gonzalez-Diaz et al., 2011a) cup products on cubical complexes have been applied to compute cohomology rings of 3D digital images (represented as cubical complexes $Q(I)$), thus simplify the combinatorial structure of $Q(I)$ and obtain a homeomorphic cellular complex $P(I)$ with fewer cells, which improves the computational efficiency. Gonzalez-Diaz et al. (2011a) think of the cup product as the way the $n$–dimensional holes obtained in homology are related to each other. It is known that two objects with non isomorphic cohomology rings are not homotopic.
Formulas for a diagonal approximation on a general polygon that are used to compute cup products on the cohomology of polygonal cellular complexes are presented in (Gonzalez-Diaz et al., 2011b). Their cup product uses a complex structure called the AT–model (Gonzalez-Diaz et al., 2011b, Theorem 2), and the coefficient of their cochains lie in the field $\mathbb{Z}_{2}$. Whereas the coefficients of our cochains lie in $\mathbb{R}$.
In common to previous approaches, our notion of discrete wedge product is metric–independent and satisfies core properties such as the Leibniz product rule, skew–commutativity, and associativity on closed forms.
2 Results
We present a new cup product on polygonal 2–pseudomanifolds (Definition 5.4) and show that previously studied cup products on simplicial and cubical 2–pseudomanifolds are a special case of this more general polygonal cup product. We proceed by proving that the polygonal cup product satisfies the Leibniz product rule with discrete exterior derivative (Lemma 5.1) and it satisfies defining properties of an abstract cup product on cellular complexes (Theorem 5.1). Moreover, in Proposition 5.3 we show that the polygonal cup product shares several further properties with the wedge product of differential forms, such skew–commutativity. We thus use the terms polygonal cup product and polygonal wedge product interchangeably.
To properly define the polygonal cup product, we first review the appropriate discrete setting in Section 3, concretely we define orientable pseudomanifolds, which is the domain of our interest, and we revise the boundary operator of chains of cells. In Section 4 we formally introduce discrete differential forms as cochains and discrete exterior derivative as coboundary operator. Section 5 then contains our the actual results announced in the previous paragraph.
3 Discrete setting
The discrete domain over which our discrete wedge product operates is a general polygonal mesh, i.e., a surface mesh whose faces are arbitrary simple polygons. From the perspective of algebraic topology such a mesh corresponds to a 2–dimensional CW complex with some additional structures: a 2–dimensional pseudomanifold. In this section we briefly revise several definitions that ultimately leads to the definition of an orientable pseudomanifold and the boundary operator.
A CW complex is a cellular complex, invented by J. H. C. Whitehead, and it is a more general complex that the simplicial one. According to (Munkres, 1984, Paragraph 38), cellular complexes are more effective for computations of homology groups of given spaces. The following set of definitions were adapted from definitions in (Munkres, 1984, Paragraph 38).
Definition 3.1.
A space is called a m-cell if it is homeomorphic to the unit $m$-dimensional $B^{m}$. It is called an open cell of dimension $m$ if it is homeomorphic to Int $B^{m}$. The set $\dot{e}_{i}^{m}:=\bar{e}_{i}^{m}-e_{i}^{m}$ for $m>0$ is called the boundary of the $m$-cell $e_{i}^{m}$.
Definition 3.2.
A CW complex is a space $X$ and a collection of disjoint open cells $e_{i}$ whose union is $X$ such that:
  1. 1.
    $X$ is Hausdorff.
  2. 2.
[…]
  3. 3.
    A set $A$ is closed in $X$ if $A\cap\bar{e}_{i}$ is closed in $\bar{e}_{i}$ for each $i$.
If the collection of open cells is finite and the maps $f_{i}$ can be taken to be homeomorphisms, and each set $\dot{e}_{i}:=\bar{e}_{i}-e_{i}$ equals the union of some open cells of $X$, then $X$ is called a finite regular CW complex.
The condition (2) from the above definition is called “closure-finiteness” and condition (3) expresses the “weak topology” relative to the collection $\{\bar{e}_{i}\}$. These terms originate the letters C and W in the expression “CW complex”.
Next, we define pseudomanifold and introduce its “orientability” through incidence numbers, which will also allow to define boundary homomorphism later on. For more details, see (Massey, 1991, Chapter IX).
Definition 3.3.
A CW n-pseudomanifold is an $n$-dimensional finite regular CW complex which satisfies the following three conditions:
  1. 1.
    Every cell is a face of some $n$-cell.
  2. 2.
    Every $(n-1)$-cell is a face of at most two $n$-cells.
  3. 3.
    Given any two $n$-cells, $e^{n}_{a}$ and $e^{n}_{b}$, there exist a sequence of $n$-cells
    $e_{0}^{n},e_{1}^{n},\dots,e^{n}_{k}$
    such that $e^{n}_{a}=e_{0}^{n},\,e^{n}_{b}=e^{n}_{k}$, and $e^{n}_{i-1}$ and $e^{n}_{i}$ have a common $(n-1)$-dimensional face ($i=1,\dots,k$).
If in the condition (2) of above definition, every $(n-1)$-dimensional cell is a face of exactly two $n$-cells, then we talk about pseudomanifold without boundary ((Massey, 1991, Chapter IX)). On the other hand, if there exist at least one $(n-1)$-cell which is incident with only one $n$-cell, we talk about pseudomanifold with boundary.
Definition 3.4.
Let $K=\{K^{n}\}$ be a pseudomanifold on the topological space $X$. And let $[e_{\lambda}^{n}:e_{\mu}^{n-1}]$ be the incidence number of the cells $e_{\lambda}^{n}$ and $e_{\mu}^{n-1}$, for $n>0$, such that:
  1. 1.
    If $e_{\mu}^{n-1}$ is not a face of $e_{\lambda}^{n}$, then $[e_{\lambda}^{n}:e_{\mu}^{n-1}]=0$.
  2. 2.
    If $e_{\mu}^{n-1}$ is a face of $e_{\lambda}^{n}$, then $[e_{\lambda}^{n}:e_{\mu}^{n-1}]=\pm 1$.
  3. 3.
    If $e_{\alpha}^{0}$ and $e_{\beta}^{0}$ are the two vertices of the 1-cell $e_{\lambda}^{1}$, then $[e_{\lambda}^{1}:e_{\alpha}^{0}]+[e_{\lambda}^{1}:e_{\beta}^{0}]=0$.
  4. 4.
    Let $e_{\lambda}^{n}$ and $e_{\rho}^{n-2}$ be cells such that $e_{\rho}^{n-2}\prec e_{\lambda}^{n}$; let $e_{\alpha}^{n-1}$ and $e_{\beta}^{n-1}$ denote the unique $(n-1)$-cells $e^{n-1}$ such that $e_{\rho}^{n-2}\prec e^{n-1}\prec e_{\lambda}^{n-1}$. Then
    $[e_{\lambda}^{n}:e_{\alpha}^{n-1}][e_{\alpha}^{n-1}:e_{\rho}^{n-2}]+[e_{\lambda}^{n}:e_{\beta}^{n-1}][e_{\beta}^{n-1}:e_{\rho}^{n-2}]=0.$
With these conditions it is possible to choose an orientation for each cell $e_{\lambda}^{n}$ in one and only one way.
Thus we can specify orientations for the cells by specifying a set of incidence numbers for the complex. Before we define boundary homomorphism, we introduce the notion of orientability of a pseudomanifold.
Definition 3.5.
Let $K$ be an $n$-dimensional pseudomanifold (possibly with or without boundary), and let $e_{1}^{n},e_{2}^{n}$ be $n$-cells with a common $(n-1)$-cell $e^{n-1}$. We define orientations for $e_{1}^{n}$ and $e_{2}^{n}$ to be coherent (with respect to $e^{n-1}$) if:
$[e_{1}^{n}:e^{n-1}]+[e_{2}^{n}:e^{n-1}]=0.$
A set of orientations for all the $n$-cells of $K$ is said to be coherent if it is coherent in the above sense for any pair of $n$-cells with a common $(n-1)$-face.
$K$ is said to be orientable if all its $n$-cells can be oriented such that any pair of $n$-cells sharing an $(n-1)$-dimensional face are oriented coherently. Otherwise it is called nonorientable.
The orientability or nonorientability of an $n$-dimensional pseudomanifold $K$ only depends on the underlying topological space involved, and not on the choice of the regular CW complex $K$.
Before we define the boundary operator, we need one more definition:
Definition 3.6.
Let $K$ be a pseudomanifold. A p-chain on $K$ is a function $c$ from the set of oriented $p$-cells of $K$ to the integers, such that $c(e)=-c({e}^{\prime})$ if $e$ and ${e^{\prime}}$ are opposite orientations of the same cell. We add $p$-chains by adding their values, the resulting group is denoted $C_{p}(K)$ and is called the chain group of $K$. If $p<0$ or $p>\dim K$, we let $C_{p}(K)$ denote the trivial group.
The incidence numbers, or equivalently, the set of orientations of $p$-cells of $K$, give us such a $p$-chain.
Definition 3.7.
Let $C_{n}(K),n\geq 0,$ be the chain groups of $K$. The boundary homomorphism
$\partial_{n}:C_{n}(K)\rightarrow C_{n-1}(K),\;n>0,$
is defined to be
$\partial_{n}(e^{n})=\sum_{\lambda}[e^{n}:e^{n-1}_{\lambda}]e^{n-1}_{\lambda},$ (1)
where $e^{n}$ is an oriented $n$-cell of $K$ with $n>0$.
From the definition of incidence numbers it follows that $\partial_{n}$ is well-defined and that $\partial_{n}(-e^{n})=-\partial_{n}(e^{n})$. The boundary homomorphism on 2–chains is illustrated in Figure 1.
Definition 3.8.
A chain complex $K=\{K_{i},\partial_{i}\}$ is a sequence of abelian groups $K_{i},i\in\mathbb{Z}$, and a sequence of homomorphisms $\partial_{i}:K_{i}\rightarrow K_{i-1}$ which are required to satisfy the condition
$\partial_{i-1}\partial_{i}=0\;\;\forall i.$
For any such chain complex $K=\{K_{i},\partial_{i}\}$ we have that $\operatorname{im}\partial_{i+1}\subset\ker\partial_{i}\subset K_{i}$ and we can define
$H_{i}(K)=\frac{\ker\partial_{i}}{\operatorname{im}\partial_{i+1}},$
called the $i$th homology group of $K$.
If $K$ is an $n$-dimensional pseudomanifold, then $K^{p}=\emptyset$ for $p<0$ and $p>n$, thus $C_{n+1}(K)=0$ and the chain complex reads:
$0\longrightarrow C_{n}(K)\overset{\partial_{n}}{\longrightarrow}\cdots\overset{\partial_{k+1}}{\longrightarrow}C_{k}(K)\overset{\partial_{k}}{\longrightarrow}\cdots\overset{\partial_{1}}{\longrightarrow}C_{0}(K)\longrightarrow 0$
4 Discrete Differential Forms and the Exterior Derivative
Having learned about chains and the boundary homomorphism we can now define cochains and the coboundary operator on them, which correspond to our discrete differential forms and the discrete exterior derivative.
Definition 4.1.
Let $K$ be a pseudomanifold and $C_{n}(K)$ the group of oriented $n$-chains of $K$. Let $G$ be an abelian group. The group of n-dimensional cochains of $K$, with coefficients in $G$, is the group
$C^{n}(K)=\text{ Hom }(C_{n}(K),G).$
The coboundary operator $d$ is defined to be the dual of the boundary operator $\partial_{n}:C_{n+1}(K)\rightarrow C_{n}(K)$, i.e., it is the homomorphism
$d:C^{n}(K)\rightarrow C^{n+1}(K),$
such that $dd=0.$
We can think of a cochain group $C^{n}(K)$ as dual of a chain group $C_{n}(K)$. We now give the formal definition of discrete differential forms and the discrete exterior derivative:
Definition 4.2.
A discrete q–form $\alpha^{q}$ on a pseudomanifold $K$ is an element of $C^{q}(K)$, the group of q–dimensional cochains of $K$, that is
$\alpha^{q}\in C^{q}(K)=\text{ Hom }(C_{q}(K),G).$
The discrete exterior derivative $d_{q}:C^{q}(K)\rightarrow C^{q+1}(K)$ is the coboundary operator and it holds:
$(d_{q}\alpha^{q})(c_{q+1})=\alpha(\partial_{q+1}c_{q+1})=\sum_{c\in C_{q}(K)}[c_{q+1}:c]\alpha(c).$ (2)
The discrete exterior derivative is illustrated in Figure 1.
Figure 1: Boundary homomorphism and exterior derivative on polygonal complex. In this figure, our pseudomanifold $K$ has 2–chain $C_{2}=(f_{0},f_{1})$, 1–chain $C_{1}=(e_{0},\dots,e_{5}$), and 0–chain $C_{0}=(v_{0},\dots,v_{4})$. If 2–form $\gamma^{2}$ was given as the exterior derivative of 1–form $\beta^{1}$, then $\gamma=d\beta$ would be defined per faces $f_{0},f_{1}$ by equation (2). Concretely $\gamma(f_{1})=d\beta(f_{1})$ is:
$(d\beta)(f_{1})=\beta(\partial f_{1})=\sum_{e_{i}\in C_{1}}[f_{1}:e_{i}]\beta(e_{i})=0\cdot\beta(e_{0})-1\cdot\beta(e_{1})+0\cdot\beta(e_{2})+1\cdot\beta(e_{3})+1\cdot\beta(e_{4})+1\cdot\beta(e_{5}).$
Since now on, we work with coefficient group $G=(\mathbb{R},+)$, which corresponds to working with real–valued discrete differential forms.
Similarly to the definition of chain complexes (Definition 3.8), we have the following:
Definition 4.3.
A cochain complex $(C^{\ast}(K,\mathbb{R}),d)$ is a sequence of abelian groups of $n$-dimensional cochains $C^{n}(K)$ together with the coboundary operators $d_{n}:C^{n}(K)\rightarrow C^{n+1}(K)$, i.e.,
$0\overset{d_{n}}{\longleftarrow}C^{n}(K)\overset{d_{n-1}}{\longleftarrow}\cdots\overset{d_{k}}{\longleftarrow}C^{k}(K)\overset{d_{k-1}}{\longleftarrow}\cdots\overset{d_{0}}{\longleftarrow}C^{0}(K)\longleftarrow 0$
A $q$–form $\alpha$ is said to be a closed form (cocycle) if $d\alpha=0$, and it is called an exact form (coboundary) if there exist a $(q-1)$–form $\beta$ such that $d\beta=\alpha$.
Note that each exact $q$-form $\alpha=d\beta$ is closed, since $d\alpha=dd\beta=0$. Two closed $q$-forms are cohomologous if they differ by an exact $q$–form. We denote cohomology classes by $H^{\ast}$:
$H^{q}(C^{\ast})=\frac{\ker d_{q}(C^{\ast})}{\operatorname{im}d_{q-1}(C^{\ast})}.$
5 Wedge and Cup Products
In this section we first present the wedge product on a manifold and then give a general definition of an abstract cup product of cellular cochains to show the clear analogy between these two. We then proceed with equations for cup products on triangles and quadrilaterals, and finally present simple formulas for cup products on general polygons (Subsection 5.2).
Wedge product
A differential $k$-form $\alpha$ on a smooth $n$–manifold $M$ is a tensor field of type $(0,k)$ that is completely antisymmetric, i.e., for $\alpha\in\Omega^{k}(M)\subset T^{0}_{k}(M)$ we have $\alpha(v,w)=-\alpha(w,v)$ for all $v,w\in TM$. Let $\{e_{1},\dots,e_{n}\}$ be a basis of the tangent space $T_{x}M$ to $M$ at point $x\in M$. Let $\alpha\in T^{0}_{k}(M)$ and $\beta\in T^{0}_{l}(M)$, we define their wedge product $\alpha\wedge\beta\in T^{0}_{k+l}(M)$ at the point $x\in M$ by
$(\alpha\wedge\beta)=\sum_{{\tau}}\operatorname{sign}{\tau}\alpha(e_{\tau(1)},\dots,e_{\tau(k)})\beta(e_{\tau(k+1)},\dots,e_{\tau(k+l)}),$ (3)
where the sum is over all permutations ${\tau}$ of $\{1,\dots,k+l\}$ such that $\tau(1)<\dots<\tau(k)$ and $\tau(k+1)<\dots<\tau(k+l)$. And we have that $\operatorname{sign}{\tau}=+1$ if the permutation is even and $\operatorname{sign}{\tau}=-1$ if it is odd.
The wedge product exhibits the following properties, which we maintain also in the discrete setting except the associativity (compare with the later Proposition 5.3):
Proposition 5.1.
For $\alpha\in T^{0}_{k}(M)$, $\beta\in T^{0}_{l}(M)$, and $\gamma\in T^{0}_{m}(M)$, we have
  1. 1.
    $\wedge$ is bilinear: $\alpha\wedge(c_{1}\beta+c_{2}\gamma)=c_{1}(\alpha\wedge\beta)+c_{2}(\alpha\wedge\gamma)$ and $(c_{1}\alpha+c_{2}\beta)\wedge\gamma=c_{1}(\alpha\wedge\gamma)+c_{2}(\beta\wedge\gamma)$ for some constants $c_{1},c_{2}\in\mathbb{R}$,
  2. 2.
    $\wedge$ is skew commutative: $\alpha\wedge\beta=(-1)^{kl}\beta\wedge\alpha$,
  3. 3.
    $\wedge$ is associative: $\alpha\wedge(\beta\wedge\gamma)=(\alpha\wedge\beta)\wedge\gamma$,
  4. 4.
    Leibniz rule: $d(\alpha\wedge\beta)=d\alpha\wedge\beta+(-1)^{k}\alpha\wedge d\beta$.
For the proof see (Abraham et al., 1988, Proposition 7.1.5 and Theorem 7.4.1).
Cup product
The wedge product allows for building higher degree forms from lower degree ones, similarly a cup product is a product of cochains of arbitrary degree $p$ and $q$ that returns a cochain of degree $p+q$.
Whitney in (Whitney, 1957, §9 of Appendix II) gives the following abstract definition of a cup product, which also appears in (Arnold, 2012, Definition 2.3.3).
Definition 5.1.
Let $X$ be a cell complex, the cup product of two cochains $c^{p}$ and $c^{q}$ is a bilinear operation $\cup:C^{p}(X)\times C^{q}(X)\rightarrow C^{p+q}(X)$ that satisfies the following three properties:
  1. 1.
    Let $\sigma_{p}\in C_{p}(X)$ and $\sigma_{q}\in C_{q}(X)$. Then $\sigma^{p}\cup\sigma^{q}$ is a $(p+q)$-cochain in $St(\sigma_{p})\cdot St(\sigma_{q})$, where $St(\sigma_{i})$ is the union of all cells in which $c_{i}$ is a face, and $A\cdot B$ denotes the union of all cells in $A$ and $B$.
  2. 2.
    $d(c^{p}\cup c^{q})=dc^{p}\cup c^{q}+(-1)^{p}c^{p}\cup dc^{q}$ (Leibniz rule).
  3. 3.
    If $X$ is connected, then there exist a real number $\gamma_{\cup}$ such that $I^{0}\cup c^{p}=c^{p}\cup I^{0}=\gamma_{\cup}c^{p}$, where $I^{0}$ is the constant 0-cochain that takes value 1 on the 0-cells of $X$.
The interested reader is also referred to (Massey, 1991, Chapter XIII). Whitney further asserts that thanks to the Leibniz rule being valid, we have the following:
Proposition 5.2.
For $\alpha\in H^{k}(X)$, $\beta\in H^{l}(X)$, $\gamma\in H^{m}(X)$, the cup product defines a bilinear operation $H^{k}(X)\times H^{l}(X)\rightarrow H^{k+l}(X)$, which is uniquely determined (well–defined) and is
  1. 1.
    skew commutative: $\alpha\cup\beta=(-1)^{kl}\beta\cup\alpha$,
  2. 2.
    associative: $\alpha\cup(\beta\cup\gamma)=(\alpha\cup\beta)\cup\gamma$.
Thus the cup endows the cellular cohomology $H^{*}(X)$ with a graded commutative ring structure, with the cohomology class of $I^{0}$ as unit element.
Comparing Proposition 5.1 with Definition 5.1 and Proposition 5.2, we affirm that both the wedge and the cup product, respectively, are bilinear operations that take as input two forms, resp. cochains, of arbitrary degree $p$ and $q$ and return a form, resp. cochain, of degree $p+q$. Moreover, they both satisfy the Leibniz rule. But a general cup products is guaranteed to be associative and skew–commutative only on cohomology. Fortunately, for our cup product on polygons we will recuperate the skew–commutativity for any discrete forms (not just closed ones). However, our polygonal cup product will be associative only on closed forms, see the Proposition 5.3.
Even though explicit formulas for computing cup product on a general cell complex are unknown, there are well–known explicit formulas for cup products on simplicial and cubical complexes (Arnold, 2012). We expand the set of known cup products with cup products on general polygons with coefficients in $\mathbb{R}$, which is one of our contributions.
5.1 Simplicial and cubical cup products
Let us first recall the definition of simplicial Whitney forms. Let $c=(v_{0},\dots,v_{p})$ be a $p$–simplex with barycentric coordinates $x_{0},\dots,x_{p}$. Thus $0\leq x_{j}\leq 1\;\forall j$ and $\sum_{i=0}^{p}=1$. The simplicial Whitney $p$–form corresponding to the simplex $c$ is
$W(c)=p!\sum_{i=0}^{p}(-1)^{i}x_{i}dx_{0}\wedge\dots\wedge\widehat{dx_{i}}\wedge\dots\wedge dx_{p},$
where $\;{}\widehat{}\;$ indicates a term omitted from the product. This definition extends linearly to define Whitney forms on all of $C^{\ast}(K)$.
Arnold in (Arnold, 2012, Definition 4.2.1) presents the following definition of a cup product on simplices, which she calls the Wilson’s cup product, $\cup:C^{p}(K)\times C^{q}(K)\rightarrow C^{p+q}(K)$ given by
$\alpha\cup\beta(c)=\int_{c}W\alpha\wedge W\beta,$ (4)
where $c$ is an $(p+q)$–simplex and $W$ is the simplicial Whitney form.
Even though the above cup product has appeared already in (Whitney, 1957), it has been studied by Scott Wilson in (Wilson, 2007), where he proved that the cup product converges to wedge product of differential forms (Wilson, 2007, Theorem 5.4). Furthermore, Arnold claims that Wilson’s work was the motivation for the majority of results presented in (Arnold, 2012), including the definition of cubical cup product presented in the next paragraphs.
Using the equation (4) we compute cup products on 2–dimensional simplicial complexes and state them as the following definition:
Definition 5.2.
Let $(v_{0},\dots,v_{k})$ be a $k$–simplex of a pseudomanifold $K$, the simplicial cup product $\cup:C^{p}(K,\mathbb{R})\times C^{q}(K,\mathbb{R})\rightarrow C^{p+q}(K,\mathbb{R})$, i.e., the cup product defined over each simplex $(v_{0},\dots,v_{k})$ reads:
$(\alpha^{0}\cup\beta^{0})(v_{0})$ $=$ $\alpha(v_{0})\beta(v_{0}),$
$(\alpha^{0}\cup\beta^{1})(v_{0},v_{1})$ $=$ $\frac{1}{2}(\alpha(v_{0})+\alpha(v_{1}))\beta(v_{0},v_{1}),$
$(\alpha^{0}\cup\beta^{2})(v_{0},v_{1},v_{2})$ $=$ $\frac{1}{3}(\alpha(v_{0})+\alpha(v_{1})+\alpha(v_{2}))\beta(v_{0},v_{1},v_{2}),$
$(\alpha^{1}\cup\beta^{1})(v_{0},v_{1},v_{2})$ $=$ $\frac{1}{6}\sum_{i=0}^{2}\alpha(v_{i},v_{i+1})(\beta(v_{i+1},v_{i+2})-\beta(v_{i-1},v_{i})),$
where the indices are taken to be modulo 3.
The cup product on 1–cochains (discrete 1–forms) is illustrated in Figure 2.
Remark.
In our derivation we have actually obtained $(\alpha^{0}\cup\beta^{2})=\frac{1}{6}(\alpha(v_{0})+\alpha(v_{1})+\alpha(v_{2}))\beta(v_{0},v_{1},v_{2})$ and not $\frac{1}{3}(\alpha(v_{0})+\alpha(v_{1})+\alpha(v_{2}))\beta(v_{0},v_{1},v_{2})$. But in the case of $\frac{1}{6}$, the Leibniz rule for $d(\alpha^{0}\cup\beta^{1})$ is only satisfied if $d\beta=0$, i.e., $\beta$ is a closed form. Whereas for $\frac{1}{3}$, the Leibniz rule for $d(\alpha^{0}\cup\beta^{1})=d\alpha\cup\beta+\alpha\cup d\beta$ is satisfied for any $\alpha$ and $\beta$.
Figure 2: The cup product of two 1–forms is a 2–form located on faces.
The simplicial cup product satisfies the properties of Proposition 5.3, which we prove therein. Comparing the Proposition 5.3 with Propositions 5.2 and 5.1, we can see that our simplicial cup product is skew–commutative on any discrete forms, not only on cohomology as the generic cup product would require, but it is still not associative in general as the wedge product, only on cohomology.
In (Arnold, 2012, Definition 3.2.1) the author defines a cubical cup product $\cup_{c}$, for which the product of two cochains agrees with the wedge product of their cubical Whitney forms (see (Arnold, 2012, Section 3.2.2)), thus the Whitney map provides a connection of cubical cohomology with de Rham cohomology ((Arnold, 2012, Theorem 3.2.12)). From (Arnold, 2012, Definition 3.2.1) we computed the cup product on $k$–dimensional cubes for $0\leq k\leq 2$ and we present the formulas in the following definition:
Definition 5.3.
The cubical cup product $\cup:C^{p}(K,\mathbb{R})\times C^{q}(K,\mathbb{R})\rightarrow C^{p+q}(K,\mathbb{R})$ on a quadrilateral pseudomanifold $K$ is defined by its action on a $k$–dimensional cube $(v_{0},\dots,v_{2^{k}-1})$ as:
$(\alpha^{0}\cup\beta^{0})(v_{0})$ $=$ $\alpha(v_{0})\beta(v_{0}),$
$(\alpha^{0}\cup\beta^{1})(v_{0},v_{1})$ $=$ $\frac{1}{2}(\alpha(v_{0})+\alpha(v_{1}))\beta(v_{0},v_{1}),$
$(\alpha^{0}\cup\beta^{2})(v_{0},v_{1},v_{2},v_{3})$ $=$ $\frac{1}{4}\Big{(}\sum_{i=0}^{3}\alpha(v_{i})\Big{)}\beta(v_{0},v_{1},v_{2},v_{3}),$
$(\alpha^{1}\cup\beta^{1})(v_{0},v_{1},v_{2},v_{3})$ $=$ $\frac{1}{4}\sum_{i=0}^{3}\alpha(v_{i},v_{i+1})(\beta(v_{i+1},v_{i+2})-\beta(v_{i-1},v_{i})),$
where the indices are taken to be modulo 4.
5.2 The cup product on polygons
As advertised earlier, we can extend the definitions of the discrete wedge product to any $p$–polygon such that it still satisfies the defining properties of a cup product (Definition 5.1) and therefore automatically also the properties of Proposition 5.2. Moreover, we show that our polygonal cup product is a skew–commutative bilinear operation on any polygonal forms, not just on closed ones, viz. the Proposition 5.3.
Let us remember that a discrete $q$–form $\alpha^{q}$ is a $q$–cochain defined per each element of a $q$–chain $C_{q}(K)$:
$\alpha^{q}=\alpha(c_{0}),\dots,\alpha(c_{p}),\;\;c_{i}\in C_{q}(K).$
Furthermore, we sum two discrete $q$–forms $\alpha^{q},\beta^{q}$ by summing their values per each element of a $q$–chain $C_{q}(K)$ and multiply a $q$–form $\alpha^{q}$ by a scalar $t\in\mathbb{R}$ in element–wise fashion too:
$\alpha^{q}+\beta^{q}$ $=(\alpha+\beta)(c_{0}),\dots,(\alpha+\beta)(c_{p}),$ (5)
$t\cdot\alpha^{q}$ $=t\cdot\alpha(c_{0}),\dots,t\cdot\alpha(c_{p}).$ (6)
Definition 5.4.
Let $K$ be a 2–pseudomanifold whose 2–cells (faces) are polygons. The cup product $C^{k}(K,\mathbb{R})\times C^{l}(K,\mathbb{R})\rightarrow C^{k+l}(K,\mathbb{R})$ of two discrete forms $\alpha^{k},\beta^{l}$ is a $(k+l)$–form $\alpha^{k}\cup\beta^{l}$ defined on each $(k+l)$–cell of $C_{k+l}(K)$ in the following fashion:
Let $f=(v_{0},\dots,v_{p-1})$ be a $p$–polygonal face, $(v_{i},v_{j})$ an edge, and $v$ a vertex of $K$, the polygonal cup product is defined for each degree and per each $(k+l)$–cell as:
$(\alpha^{0}\cup\beta^{0})(v)$ $=\alpha(v)\beta(v),$ (7)
$(\alpha^{0}\cup\beta^{1})(v_{i},v_{j})$ $=\frac{1}{2}(\alpha(v_{i})+\alpha(v_{j}))\beta(v_{i},v_{j}),$ (8)
$(\alpha^{0}\cup\beta^{2})(f)$ $=\frac{1}{p}\bigg{(}\sum_{i=0}^{p-1}\alpha(v_{i})\bigg{)}\beta(f),$ (9)
$(\alpha^{1}\cup\beta^{1})(f)$ $=\sum_{a=1}^{\lfloor\frac{p-1}{2}\rfloor}\bigg{(}\frac{1}{2}-\frac{a}{p}\bigg{)}\sum_{i=0}^{p-1}\alpha(i)(\beta((i+a)\%p)-\beta((i-a)\%p)),$ (10)
where $\alpha(i):=\alpha(e_{i})=\alpha(v_{i},v_{(i+1)\%p})$ and %$\%p$ means modulo $p$.
It is easy to see that the simplicial and cubical cup products (Definitions 5.2 and 5.3) are just special cases of the product defined above.
In Theorem 5.1 we prove that the polygonal cup product actually is a cup product, that is, we prove that it satisfies the defining properties of a cup product (Definition 5.1). But we first prove, in the following lemma, that it obeys the Leibniz rule with the discrete derivative.
Lemma 5.1.
The discrete wedge product from Definition 5.4 satisfies the Leibniz rule with the discrete derivative, that is:
$d_{0}(\alpha^{0}\cup\beta^{0})$ $=d_{0}\alpha\cup\beta+\alpha\cup d_{0}\beta,$
$d_{1}(\alpha^{0}\cup\beta^{1})$ $=d_{0}\alpha\cup\beta+\alpha\cup d_{1}\beta.$
For $d(\alpha^{0}\cup\beta^{2})$ and $d(\alpha^{1}\cup\beta^{1})$ we get trivially 0 because there are no 3–dimensional chains (and thus no 3–forms) on a 2–dimensional pseudomanifold.
Proof.
We will prove the Leibniz rule for each degree separately. The case $d(\alpha^{0}\cup\beta^{0})$: let $e=(v_{i},v_{j})$ be an edge of $K$, then
$(d_{0}\alpha\cup\beta+\alpha\cup d_{0}\beta)(e)$ $=(\alpha(v_{j})-\alpha(v_{i}))\frac{\beta(v_{i})+\beta(v_{j})}{2}$
$+\frac{\alpha(v_{i})+\alpha(v_{j})}{2}(\beta(v_{j})-\beta(v_{i}))$
$=\frac{\beta(v_{i})(-2\alpha(v_{i}))+\beta(v_{j})(2\alpha(v_{j}))}{2}$
$=\alpha(v_{j})\beta(v_{j})-\alpha(v_{i})\beta(v_{i})$
$=(\alpha\cup\beta)(v_{j})-(\alpha\cup\beta)(v_{i})$
$=d_{0}(\alpha^{0}\cup\beta^{0})(e).$
The case $d(\alpha^{0}\cup\beta^{1})$: Let us first simplify the notation and denote $\alpha_{i}=\alpha^{0}(v_{i})$, $\beta_{i}=\beta^{1}(e_{i})$. Let also any index $i$ to be understood as $i\%n$. Let $f=(v_{0},\dots,v_{n-1})$ be an $n$–polygonal face with edges $e_{i}=(v_{i},v_{i+1})$. Due to the indexes being cyclic (modulo $n$), we can reorder the following sums in this manner:
$d(\alpha^{0}\cup\beta^{1})(f)$ $=\sum_{i=0}^{n-1}\frac{\alpha_{i}+\alpha_{i+1}}{2}\beta_{i}=\sum_{i=0}^{n-1}\frac{\beta_{i-1}+\beta_{i}}{2}\alpha_{i}$
$(d\alpha^{0}\cup\beta^{1})(f)$ $=\sum_{i=0}^{n-1}\sum_{a=1}^{\lfloor\frac{n-1}{2}\rfloor}\bigg{(}\frac{1}{2}-\frac{a}{n}\bigg{)}(\alpha_{i+1}-\alpha_{i})(\beta_{i+a}-\beta_{i-a})$
$=\sum_{i=0}^{n-1}\sum_{a=1}^{\lfloor\frac{n-1}{2}\rfloor}\frac{n-2a}{2n}(\alpha_{i})(\beta_{i+a-1}-\beta_{i-a-1}-\beta_{i+a}+\beta_{i-a})$
$(\alpha^{0}\cup d\beta^{1})(f)$ $=\frac{1}{n}\sum_{i=0}^{n-1}\alpha_{i}\sum_{i=0}^{n-1}\beta_{i}$
Thus we want to show that
$\sum_{i=0}^{n-1}\alpha_{i}\frac{\beta_{i-1}+\beta_{i}}{2}=\sum_{i=0}^{n-1}\frac{\alpha_{i}}{2n}\bigg{[}\sum_{a=1}^{\lfloor\frac{n-1}{2}\rfloor}(n-2a)(\beta_{i+a-1}-\beta_{i+a}+\beta_{i-a}-\beta_{i-a-1})+2\sum_{j=0}^{n-1}\beta_{j}\bigg{]},$
but instead we will show that the following equation (11) holds for any $i\in\{0,\dots,n-1\}$, which implies that also the equality above is true.
$\frac{\beta_{i-1}+\beta_{i}}{2}=\frac{1}{2n}\bigg{[}\sum_{a=1}^{\lfloor\frac{n-1}{2}\rfloor}(n-2a)(\beta_{i+a-1}-\beta_{i+a}+\beta_{i-a}-\beta_{i-a-1})+2\sum_{j=0}^{n-1}\beta_{j}\bigg{]}$ (11)
If $n$ is even, then $a=1,\dots,\lfloor\frac{n-1}{2}\rfloor$ is equivalent to $a=1,2,\dots,\frac{n}{2}-1$, and if $n$ is odd, then it is equivalent to $a=1,\dots,\frac{n-1}{2}$. Thus we will first show the validity of (11) for $n$ even and then for $n$ odd.
For $n$ even we compute:
$\sum_{a=1}^{\frac{n}{2}-1}(n-2a)$ $(\beta_{k+a-1}-\beta_{k+a})=(n-2)(\beta_{k}-\beta_{k+1})+(n-4)(\beta_{k+1}-\beta_{k+2})$
$\;+(n-6)(\beta_{k+2}-\beta_{k+3})+\cdots+6(\beta_{k+\frac{n}{2}-4}-\beta_{k+\frac{n}{2}-3})$
$\;+4(\beta_{k+\frac{n}{2}-3}-\beta_{k+\frac{n}{2}-2})+2(\beta_{k+\frac{n}{2}-2}-\beta_{k+\frac{n}{2}-1})$
$=$ $\;n\beta_{k}-2(\beta_{k}+\beta_{k+1}+\beta_{k+2}+\beta_{k+3}+\dots$
$+\beta_{k+\frac{n}{2}-3}+\beta_{k+\frac{n}{2}-2}+\beta_{k+\frac{n}{2}-1}),$
$\sum_{a=1}^{\frac{n}{2}-1}(n-2a)$ $(\beta_{k-a}-\beta_{k-a-1})=(n-2)(\beta_{k-1}-\beta_{k-2})+(n-4)(\beta_{k-2}-\beta_{k-3})$
$\;+(n-6)(\beta_{k-3}-\beta_{k-4})+\cdots+6(\beta_{k-\frac{n}{2}+3}-\beta_{k-\frac{n}{2}+2})$
$\;+4(\beta_{k-\frac{n}{2}+2}-\beta_{k-\frac{n}{2}+1})+2(\beta_{k-\frac{n}{2}+1}-\beta_{k-\frac{n}{2}})$
$=$ $\;n\beta_{k-1}-2(\beta_{k-1}+\beta_{k-2}+\beta_{k-3}+\dots$
$+\beta_{k+\frac{n}{2}+2}+\beta_{k+\frac{n}{2}+1}+\beta_{k+\frac{n}{2}}).$
Thus
$\sum_{a=1}^{\frac{n}{2}-1}(n-2a)(\beta_{k+a-1}-\beta_{k+a}+\beta_{k-a}-\beta_{k-a-1})$ $=n(\beta_{k-1}+\beta_{k})-2\sum_{i=k}^{k+n-1}\beta_{i}$
$=n(\beta_{k-1}+\beta_{k})-2\sum_{j=0}^{n-1}\beta_{j}$
and for $n$ even we obtain
$\frac{1}{2n}\bigg{[}\sum_{a=1}^{\lfloor\frac{n-1}{2}\rfloor}(n-2a)(\beta_{k+a-1}-\beta_{k+a}+\beta_{k-a}-\beta_{k-a-1})+2\sum_{j=0}^{n-1}\beta_{j}\bigg{]}=\frac{\beta_{k-1}+\beta_{k}}{2}.$
For $n$ odd we proceed in a similar fashion:
$\sum_{a=1}^{\frac{n-1}{2}}(n-2a)$ $(\beta_{k+a-1}-\beta_{k+a})=(n-2)(\beta_{k}-\beta_{k+1})+(n-4)(\beta_{k+1}-\beta_{k+2})$
$\;+(n-6)(\beta_{k+2}-\beta_{k+3})+\cdots+5(\beta_{k+\frac{n-1}{2}-3}-\beta_{k+\frac{n-1}{2}-2})$
$\;+3(\beta_{k+\frac{n-1}{2}-2}-\beta_{k+\frac{n-1}{2}-1})+2(\beta_{k+\frac{n-1}{2}-1}-\beta_{k+\frac{n-1}{2}})$
$=$ $\;n\beta_{k}-2(\beta_{k}+\beta_{k+1}+\beta_{k+2}+\beta_{k+3}+\dots+\beta_{k+\frac{n-1}{2}-2}+\beta_{k+\frac{n-1}{2}-1})$
$-\beta_{k+\frac{n-1}{2}},$
$\sum_{a=1}^{\frac{n-1}{2}}(n-2a)$ $(\beta_{k-a}-\beta_{k-a-1})=(n-2)(\beta_{k-1}-\beta_{k-2})+(n-4)(\beta_{k-2}-\beta_{k-3})$
$\;+(n-6)(\beta_{k-3}-\beta_{k-4})+\cdots+5(\beta_{k-\frac{n-1}{2}+2}-\beta_{k-\frac{n-1}{2}+1})$
$\;+3(\beta_{k-\frac{n-1}{2}+1}-\beta_{k-\frac{n-1}{2}})+(\beta_{k-\frac{n-1}{2}}-\beta_{k-\frac{n-1}{2}-1})$
$=$ $\;n\beta_{k-1}-2(\beta_{k+\frac{n-1}{2}+1}+\beta_{k+\frac{n-1}{2}+2}+\dots+\beta_{k-3}+\beta_{k-2}+\beta_{k-1})$
$-\beta_{k+\frac{n-1}{2}}.$
Thus
$\sum_{a=1}^{\frac{n-1}{2}}(n-2a)(\beta_{k+a-1}-\beta_{k+a}+\beta_{k-a}-\beta_{k-a-1})$ $=n(\beta_{k-1}+\beta_{k})-2\sum_{i=k}^{k+n-1}\beta_{i}$
$=n(\beta_{k-1}+\beta_{k})-2\sum_{j=0}^{n-1}\beta_{j}$
and for $n$ odd we get
$\frac{1}{2n}\bigg{[}\sum_{a=1}^{\lfloor\frac{n-1}{2}\rfloor}(n-2a)(\beta_{k+a-1}-\beta_{k+a}+\beta_{k-a}-\beta_{k-a-1})+2\sum_{j=0}^{n-1}\beta_{j}\bigg{]}=\frac{\beta_{k-1}+\beta_{k}}{2}.$
We have shown that the equality (11) holds for any $i\in\{0,1,\dots,n-1\}$, which implies $d_{1}(\alpha^{0}\cup\beta^{1})=d_{0}\alpha\cup\beta+\alpha\cup d_{1}\beta$. ∎
Now we are ready to claim that our discrete wedge product on polygonal pseudomanifolds is a cup product:
Theorem 5.1.
The polygonal cup product from Definition 5.4 is a cup product, i.e., it satisfies the Definition 5.1.
Proof.
From the formulas (7)–(10) and the way we sum two discrete forms (eq. 5) or multiply them by scalar (eq. 6), it is a simple exercise to see that $\forall t\in\mathbb{R},\;\forall c_{i}\in C_{k+l},0\leq(k+l)\leq 2$:
$(t\cdot\alpha^{k}\cup\beta^{l})(c_{i})$ $=(\alpha^{k}\cup t\cdot\beta^{l})(c_{i}),$
$((\alpha^{k}_{1}+\alpha^{k}_{2})\cup\beta^{l})(c_{i})$ $=(\alpha^{k}_{1}\cup\beta^{l})(c_{i})+(\alpha^{k}_{2}\cup\beta^{l})(c_{i}),$
$(\alpha^{k}\cup(\beta^{l}_{1}+\beta^{l}_{2})(c_{i})$ $=(\alpha^{k}\cup\beta^{l}_{1})(c_{i})+(\alpha^{k}\cup\beta^{l}_{2})(c_{i}).$
Therefore the polygonal cup product is a bilinear operation.
It also satisfies item 1 of Definition 5.1 – the result of product of two forms is located on a face that is incident to both the elements where the two original discrete forms reside.
Due to Lemma 5.1, our product satisfies item 2, the Leibniz rule.
We will now prove that item 3 of Definition 5.1 also holds. A pseudomanifold $K$ as we defined it is a connected cell complex, therefore we have to show that there exist a real number $\xi$ such that $I^{0}\cup\alpha^{p}=\alpha^{p}\cup I^{0}=\xi\alpha^{p}$, where $I^{0}$ is the constant 0–form that takes value 1 on the vertices of $K$. Using the equations of Definition 5.4, we can thus write:
$I^{0}\cup\alpha^{0}$ $=\sum_{v_{i}\in V(K)}I^{0}\cup\alpha^{0}(v_{i})=\sum_{v_{i}\in V}I^{0}(v_{i})\alpha^{0}(v_{i})=\sum_{v_{i}\in V}1\alpha^{0}(v_{i})=1\alpha^{0},$
$\alpha^{0}\cup I^{0}$ $=\sum_{v_{i}\in V}\alpha^{0}\cup I^{0}(v_{i})=\sum_{v_{i}\in V}\alpha^{0}(v_{i})I^{0}(v_{i})=\sum_{v_{i}\in V}\alpha^{0}(v_{i})1=1\alpha^{0},$
$I^{0}\cup\alpha^{1}$ $=\sum_{(v_{i},v_{j})\in E(K)}I^{0}\cup\alpha^{1}(v_{i},v_{j})=\sum_{(v_{i},v_{j})\in E}\frac{I^{0}(v_{i})+I^{0}(v_{j})}{2}\alpha^{1}(v_{i},v_{j})$
$=\sum_{(v_{i},v_{j})\in E}\frac{1+1}{2}\alpha^{1}(v_{i},v_{j})=1\alpha^{1},$
$\alpha^{1}\cup I^{0}$ $=\sum_{(v_{i},v_{j})\in E}\alpha^{1}\cup I^{0}(v_{i},v_{j})=\sum_{(v_{i},v_{j})\in E}\alpha^{1}(v_{i},v_{j})\frac{I^{0}(v_{i})+I^{0}(v_{j})}{2}$
$=\sum_{(v_{i},v_{j})\in E}\alpha^{1}(v_{i},v_{j})\frac{1+1}{2}=1\alpha^{1},$
$I^{0}\cup\alpha^{2}$ $=\sum_{f_{i}\in F(K)}I^{0}\cup\alpha^{2}(f_{i})=\sum_{f_{i}\in F}\sum_{v_{j}\prec f_{i}}\frac{I^{0}(v_{j})}{p_{i}}\alpha^{2}(f_{i})$
$=\sum_{f_{i}\in F}\frac{p_{i}}{p_{i}}\alpha^{2}(f_{i})=1\alpha^{2},$
$\alpha^{2}\cup I^{0}$ $=\sum_{f_{i}\in F}\alpha^{2}\cup I^{0}(f_{i})=\sum_{f_{i}\in F}\alpha^{2}(f_{i})\sum_{v_{j}\prec f_{i}}\frac{I^{0}(v_{j})}{p_{i}}$
$=\sum_{f_{i}\in F}\alpha^{2}(f_{1})\frac{p_{i}}{p_{i}}=1\alpha^{2},$
where $p_{i}$ is the number of vertices of a face $f_{i}$. We thus conclude that $\xi=1$ satisfies $I^{0}\cup\alpha^{p}=\alpha^{p}\cup I^{0}=\xi\alpha^{p}\;\;\forall p\in\{0,1,2\}$. ∎
In the following proposition, which can be thought of as an analog of Proposition 5.1, we summarize and prove further properties of our polygonal cup product.
Proposition 5.3.
Let $K$ be a 2–pseudomanifold, $\alpha^{k}\in C^{k}(K)$, $\beta^{l}\in C^{l}(K)$, and $\gamma^{m}\in C^{m}(K)$. The polygonal cup product from Definition 5.4 satisfies these properties:
  1. 1.
    Bilinearity.
  2. 2.
    Skew–commutativity $\alpha^{k}\cup\beta^{l}=(-1)^{kl}\beta^{l}\cup\alpha^{k}$.
  3. 3.
    Leibniz rule.
  4. 4.
    Associativity under the following conditions:
    $\alpha^{0}\cup(\beta^{0}\cup\gamma^{0})=(\alpha^{0}\cup\beta^{0})\cup\gamma^{0}$ always,
    $\alpha^{0}\cup(\beta^{0}\cup\gamma^{1})=(\alpha^{0}\cup\beta^{0})\cup\gamma^{1}$ if $d\alpha^{0}=0$ or $d\beta^{0}=0$,
    $\alpha^{0}\cup(\beta^{0}\cup\gamma^{2})=(\alpha^{0}\cup\beta^{0})\cup\gamma^{2}$ if $d\alpha^{0}=0$ or $d\beta^{0}=0$,
    $\alpha^{0}\cup(\beta^{1}\cup\gamma^{1})=(\alpha^{0}\cup\beta^{1})\cup\gamma^{1}$ if $d\alpha^{0}=0$.
Proof.
From Theorem 5.1 we know our product is a bilinear operation satisfying the Leibniz rule with discrete exterior derivative. We now show that it is also skew–commutative on any discrete forms and associative in the declared fashion.
With respect to the announced form of associativity, in the case of the product of three 0–forms we get to multiplying three real numbers on each vertex, which is indeed an associative operation. For the other degrees, we have learned in the proof of Theorem 5.1 that if $K$ is connected and a 0–form $\alpha^{0}$ is constant (which is equivalent to closed), that is $\alpha^{0}(v_{i})=\rho\in\mathbb{R}\;\forall v_{i}\in K$, then the cup product of $\alpha^{0}$ with any $q$–form $\omega^{q}$ is equal to multiplying $\omega^{q}$ by the real number $\rho$. This fact yields:
$(\alpha^{0}\cup\beta^{k})\cup\gamma^{l}=(\rho\beta^{k})\cup\gamma^{l}=\rho(\beta^{k}\cup\gamma^{l})=\alpha^{0}\cup(\beta^{k}\cup\gamma^{l}),$
which proves the statement of item 4.
Concerning the skew–commutativity, looking at the equations (7)–(9) it is easy to see that $\alpha^{0}\cup\beta^{l}=\beta^{l}\cup\alpha^{0}$ for $l=0,1,2$. The case of the product of two 1–forms is slightly more involving. We need to prove that $(\alpha^{1}\cup\beta^{1})(f)=-(\beta^{1}\cup\alpha^{1})(f)$ on an $n$–polygonal face $f=(v_{0},\dots,v_{n-1})$. Let us remember that
$(\alpha^{1}\cup\beta^{1})(f)=\sum_{a=1}^{\lfloor\frac{n-1}{2}\rfloor}\bigg{(}\frac{1}{2}-\frac{a}{n}\bigg{)}\sum_{i=0}^{n-1}\alpha(i)(\beta((i+a)\%n)-\beta((i-a)\%n),$
where $\alpha(i):=\alpha(e_{i})=\alpha(v_{i},v_{(i+1)\%n})$ and %$\%n$ means modulo $n$. Again, to simplify the notation we will omit writing the symbol %$\%n$ in the indexes and by some index $k$ we will always understand $k\%n$. We next show that
$\sum_{i=0}^{n-1}\alpha(i)(\beta(i+a)-\beta(i-a))=-\sum_{i=0}^{n-1}\beta(i)(\alpha(i+a)-\alpha(i-a)).$
Due to the indexes being cyclic (modulo $n$), we can write:
$\sum_{i=0}^{n-1}\alpha(i)\beta(i+a)$ $=\alpha(0)\beta(a)+\alpha(1)\beta(1+a)+\cdots+\alpha(n-a-1)\beta(n-1)$
$+\alpha(-a)\beta(0)+\alpha(1-a)\beta(1)+\alpha(2-a)\beta(2)+\cdots$
$+\alpha(n-1)\beta(n-1+a)$
$=\sum_{j=0}^{n-1}\beta(j)\alpha(j-a),$
$\sum_{i=0}^{n-1}\alpha(i)\beta(i-a)$ $=\alpha(0)\beta(-a)+\alpha(1)\beta(1-a)+\cdots+\alpha(n+a-1)\beta(n-1)$
$+\alpha(a)\beta(0)+\alpha(1+a)\beta(1)+\alpha(2+a)\beta(2)+\cdots$
$+\alpha(n-1)\beta(n-1-a)$
$=\sum_{j=0}^{n-1}\beta(j)\alpha(j+a).$
Thus
$(\alpha^{1}\cup\beta^{1})(f)$ $=-\sum_{a=1}^{\lfloor\frac{n-1}{2}\rfloor}\bigg{(}\frac{1}{2}-\frac{a}{n}\bigg{)}\sum_{i=0}^{n-1}\beta(i)\big{(}\alpha((i+a)\%n)-\alpha((i-a)\%n)\big{)}$
$=-(\beta^{1}\cup\alpha^{1})(f).$
Hence the polygonal cup product is skew–commutative on cochain level. ∎
6 Conclusion
We have presented a new definition of discrete wedge product on polygonal 2–pseudomanifolds (Definition 5.4) and proved that this polygonal wedge product satisfies the Leibniz product rule with discrete exterior derivative (Lemma 5.1). We have also shown that the polygonal wedge product is a cup product (Theorem 5.1). Moreover, in Proposition 5.3 we have proved that the polygonal wedge product shares further properties with the wedge product of differential forms, such skew–commutativity.
Declaration of competing interest
The author declares that she has no competing interests.
Acknowledgement
The author was supported in part by Conselho Nacional de Desenvolvimento Científico e Tecnológico: processos 140396/2013-6, 300735/2017-0; and in part by the Czech Science Foundation Grant 25-15714S.